Write the equation of tangent to
3x^2+4y^2-2x+6y-25=0 at point (3,2)

Answer by Alan3354(69443)

Write the equation of tangent to
3x^2+4y^2-2x+6y-25=0 at point (3,2)
-------------
3x^2+4y^2-2x+6y-25=0
3x^2+4y^2-2x+6y = 25
3x^2 - 2x + 4y^2 + 6y = 25
===================
The point is not on the ellipse.